Suppose a basketball player has made 217 out of 302 free throws. If the player makes the next 3 free throws, I will pay you $23. Otherwise you pay me $15. Step 1 of 2 : Find the expected value of the proposition. Round your answer to two decimal places. Losses must be expressed as negative values. Answer
Answer #1

P(making a free throw) = 217/302

P(he making the next 3 free throws) = (217/302)3

= 0.371

Winning amount = $23

P(winning) = 0.371

Losing amount = $15

P(losing) = 1 - 0.371 = 0.629

Expected value = 0.371x23 - 0.629x15

= $ -0.90